# CE-QArg: Counterfactual Explanations for Quantitative Bipolar Argumentation Frameworks

### v1

@@ 1 Introduction
In this scenario, the bank rejects the applicant because the final strength is lower than the desired strength ($\sigma_{DF}(\alpha)=0.165<0.5$). To explain such an outcome, for instance, the Shapley-based attribution method (?), finds that $\beta$ has a positive importance score of $0.0975$ wrt. $\alpha$, while $\gamma$, $\zeta$, and $\rho$ all have negative importance scores of $-0.34$, $-0.0525$ and $-0.04$, respectively [^1] See the Supplementary Material (SM) for more details of the computation of strengths and importance scores.. Since the negative scores outweigh the positive one, the final strength is lower than the desired strength, resulting in an unsuccessful application.
@@
The proofs of all technical results are given in the SM.
@@ 2 Related Work
Counterfactual explanations are typically used in XAI (?) to indicate paths towards “algorithmic recourse”. Counterfactual explanations for QBAFs have not been well-studied. A recent study by (?) focuses on explaining why the strengths’ partial order of two topic arguments swap after updating QBAFs (e.g. by adding/removing arguments/edges or changing the base scores of arguments), seeing counterfactual explanations as argument sets whose elements, when updated, cause this swap. For example, suppose there are two topic arguments $\alpha$ and $\beta$ in a QBAF and the final strength of $\alpha$ is smaller than that of $\beta$. If a supporter $\gamma$ for $\alpha$ were added, resulting in $\alpha$’s strength becoming larger than $\beta$’s, then the set $\{\gamma\}$ is a possible counterfactual explanation for the strength swap between $\alpha$ and $\beta$. Instead, we focus on explaining arguments’ strength rather than partial orders between arguments’ strengths. Additionally, we focus on counterfactuals as base score functions in structure-fixed QBAFs rather than as argument sets in structure-changeable QBAFs. Another work (?), implicitly relates to counterfactual explanations for QBAFs under gradual semantics, by studying the inverse problem of identifying base score functions that can lead to a desired ranking of strengths for all arguments. Differently, we focus on reaching a desired strength for topic arguments instead of a desired ranking of all arguments.
@@ 3 Preliminaries
Concretely, we will use the Quadratic Energy (QE) (?), Restricted Euler-based (REB) (?), and Discontinuity-Free Quantitative Argumentation Debates (DF-QuAD) (?): we recap them in the SM.
@@ 5.1 Algorithm Design Properties
We can restrict the update direction of arguments based on their polarity assuming that $\sigma$ respects directionality and monotonicity. Directionality (?) states that an argument can affect only arguments that can be reached from a directed path.
@@ 5.2 Algorithms
Algorithm 1 computes the polarity from $\alpha$ to $\beta$ ($\alpha\neq\beta$) in three steps. Firstly, we compute all the non-cyclic paths from $\alpha$ to $\beta$ with Depth-First Search (DFS) and store them in a set of paths called $P_{\alpha\mapsto\beta}$. If there is no path, then the polarity is neutral. Secondly, all the nodes in all the paths are checked for cycles with the function find_cycles. If any of the node is part of a cycle, then we check whether each cycle contains an odd number of attacks. If this is the case, then the polarity from $\alpha$ to $\beta$ is unknown because the cycle will contain both odd and even numbers of attacks by going through the cycle an odd or even number of times. For instance, suppose a QBAF consists of two arguments $\alpha$ and $\beta$, where $\alpha$ attacks $\beta$, and $\beta$ supports $\alpha$. In this case, path $\langle(\alpha,\beta)\rangle$ is negative while $\langle(\alpha,\beta),(\beta,\alpha),(\alpha,\beta)\rangle$ is positive, thus $\alpha$ is unknown to $\beta$. Finally, we check the number of attacks on every path: if all the paths contain an odd (even) number of attacks, $\alpha$ is negative (positive) to $\beta$, and it is unknown otherwise.
@@
Finally, Algorithm 3 (which we call CE-QArg for Counterfactual Explanations for Quantitative bipolar Argumentation frameworks) is an iterative updating algorithm for identifying valid and cost-effective $\delta$-approximate counterfactuals. CE-QArg essentially involves determining the direction and magnitude of arguments. For brevity, we assume the current strength of the topic argument is less than the desired one. Firstly, we compute the polarity and priority for each argument with the function func_polarity and func_priority from Algorithm 1 and 2, respectively. Secondly, the key part of this algorithm is to determine the updating direction in every updating iteration. For this, we need an update list to record the updating direction in each iteration. For positive, negative and neutral arguments (lines 5-10), we only need to identify the updating direction once since they remain invariant, whereas, for unknown arguments (lines 13-20), we need to compute their updating direction in every iteration by difference quotient using the function func_dquo, which can be intuitively implemented by Definition 16 (so we omit the details). Once the updating direction of every argument is determined in an iteration, we update the base scores of all arguments all in one go by a small step multiplied by their priority (line 22) and make sure they are within the bounds. We assume the step is small enough that the interactions among arguments can be neglected. We repeat this procedure iteratively until the current strength reaches the desired strength, after which we return a possible counterfactual (lines 11-24).
@@ 7 Evaluations
We report on experiments with the QE semantics [^4] We give results with DF-QuAD and REB in the SM.. We set the updating step $\varepsilon$ to 0.01 in each iteration. Besides, we used both $L_{1}$ and $L_{2}$-norm distance as the metric for cost.
@@ 7.1 Experiment 1: Effectiveness
Table 1: Ablation studies for polarity and priority on acyclic/cyclic QBAFs: Comparison of average validity, $L_{1}$, $L_{2}$-norm distance, and runtime over 100 random generated acyclic full binary tree-like QBAFs with a depth of 7 and cyclic QBAFs with 100 arguments and 100 relations. [^6] We show the effectiveness on larger QBAFs in the SM.(The best results are shown in bold.)
@@
We first discuss the results of acyclic QBAFs (left side of slash in Table 6). All methods had the best validity of $1.00$. The use of priority is expected to shorten the $L_{1}$ and $L_{2}$-norm distance by updating the arguments close to the topic argument with a larger step. The results of BL and BL+pri showed that applying priority resulted in a 75.5% and 70.3% decrease in $L_{1}$ and $L_{2}$-norm distance, respectively, which also reduced the runtime by 23.7% in the meanwhile. These findings are in line with our hypothesis in that priority enables reaching a desired strength more cost-effectively, making the distance shorter and thus requiring fewer iterations. Utilizing polarity is expected to lower the runtime by computing the polarity only once for neutral, positive, and negative arguments in a QBAF. The results of BL and BL+pol showed that applying polarity significantly decreased the runtime by 98.7% as expected because the BL computes the difference quotient in every iteration for every argument, which causes the runtime wastage. For our CEQArg, we see that the runtime is better than that of any of the previous three algorithms. Compared to BL+pri, although the $L_{1}$ and $L_{2}$-norm distance is slightly longer in our CE-QArg, the runtime is substantially decreased.
@@ 7.3 Experiment 3: Robustness
In addition, inspired by (???), we propose a robustness against noisy execution metric, which requires that the generated counterfactual could still lead to similar output (final strength for the topic argument) even if the counterfactual is perturbed. Still taking the loan application as an example, robustness against noisy execution ensures that when a rejected applicant is very close to the provided counterfactual, then the output of this counterfactual should also be close to the desired final strength. We evaluate robustness against noisy execution by the absolute difference between the strength of a topic argument obtained with two similar counterfactuals.
@@
Figure 7: Robustness evaluation of perturbing base score functions (left) and perturbing generated counterfactuals (right) over 100 random generated QBAFs with increasing perturbation $e$ from $10^{-8}$ to $10^{-1}$. Acyclic QBAFs are binary trees with a depth of 7 while cyclic QBAFs contain 100 arguments and relations. [^8] We show the robustness on larger QBAFs in the SM.
@@ Appendix A Gradual Semantics
The DF-QuAD semantics is computed as follows:
Aggregation function:
$E_{\alpha}=\prod_{\left\{\beta\in\mathcal{A}\mid(\beta,\alpha)\in\mathcal{R^{-}}\right\}}(1-\sigma^{DF}(\beta))-\prod_{\left\{\beta\in\mathcal{A}\mid(\beta,\alpha)\in\mathcal{R^{+}}\right\}}(1-\sigma^{DF}(\beta))$ (5)
Influence function:
$\sigma^{DF}(\alpha)=\begin{cases}\tau(\alpha)-\tau(\alpha)\cdot\left|E_{\alpha}\right|&if\ E_{\alpha}\leq 0\\ \tau(\alpha)+(1-\tau(\alpha))\cdot E_{\alpha}&if\ E_{\alpha}>0\\ \end{cases}$ (6)
@@ Appendix B Additional Details for Section 1
Computation of the strength: According to the Figure 1, we have $\tau(\alpha)=0.5$, $\tau(\beta)=0.3$, $\tau(\gamma)=0.6$, $\tau(\rho)=0.7$, and $\tau(\zeta)=0.4$. According to the DF-QuAD semantics in Equation 5 and 6, we have $E_{\alpha}=0.24-0.91=-0.67$, $E_{\beta}=-0.7$, $E_{\gamma}=0.4$, $E_{\rho}=0$, $E_{\zeta}=0$, $\sigma^{DF}(\alpha)=\tau(\alpha)-\tau(\alpha)\cdot\left|E_{\alpha}\right|=0.165$, $\sigma^{DF}(\beta)=\tau(\beta)-\tau(\beta)\cdot\left|E_{\beta}\right|=0.09$, $\sigma^{DF}(\gamma)=\tau(\gamma)+(1-\tau(\gamma))\cdot E_{\gamma}=0.76$, $\sigma^{DF}(\rho)=\tau(\rho)=0.7$, $\sigma^{DF}(\zeta)=\tau(\zeta)=0.4$.
@@ Appendix D Additional Results for Section 7
Please see the tables in the following pages.

### v2

@@ 1 Introduction
In this scenario, the bank rejects the applicant because the final strength is lower than the desired strength ($\sigma_{DF}(\alpha)=0.165<0.5$). To explain such an outcome, for instance, the Shapley-based attribution method (?), finds that $\beta$ has a positive importance score of $0.0975$ wrt. $\alpha$, while $\gamma$, $\zeta$, and $\rho$ all have negative importance scores of $-0.34$, $-0.0525$ and $-0.04$, respectively [^1] See Section 3 and https://arxiv.org/abs/2407.08497 for more details of the computation of strengths and importance scores, respectively.. Since the negative scores outweigh the positive one, the final strength is lower than the desired strength, resulting in an unsuccessful application.
@@
The proofs of all technical results are in https://arxiv.org/abs/2407.08497. The code is available at https://github.com/XiangYin2021/CE-QArg.
@@ 2 Related Work
Counterfactual explanations are typically used in XAI (?) and contestable AI (??) to indicate paths towards “algorithmic recourse”. Counterfactual explanations for QBAFs have not been well-studied. A recent study by (?) focuses on explaining why the strengths’ partial order of two topic arguments swap after updating QBAFs (e.g. by adding/removing arguments/edges or changing the base scores of arguments), seeing counterfactual explanations as argument sets whose elements, when updated, cause this swap. For example, suppose there are two topic arguments $\alpha$ and $\beta$ in a QBAF and the final strength of $\alpha$ is smaller than that of $\beta$. If a supporter $\gamma$ for $\alpha$ were added, resulting in $\alpha$’s strength becoming larger than $\beta$’s, then the set $\{\gamma\}$ is a possible counterfactual explanation for the strength swap between $\alpha$ and $\beta$. Instead, we focus on explaining arguments’ strength rather than partial orders between arguments’ strengths. Additionally, we focus on counterfactuals as base score functions in structure-fixed QBAFs rather than as argument sets in structure-changeable QBAFs. Another work (?), implicitly relates to counterfactual explanations for QBAFs under gradual semantics, by studying the inverse problem of identifying base score functions that can lead to a desired ranking of strengths for all arguments. Differently, we focus on reaching a desired strength for topic arguments instead of a desired ranking of all arguments.
@@ 3 Preliminaries
Concretely, we will use the Quadratic Energy (QE) (?), Restricted Euler-based (REB) (?), and Discontinuity-Free Quantitative Argumentation Debates (DF-QuAD) (?). To aid understanding, we will show the definition and an example of the DF-QuAD semantics and recap the other two in https://arxiv.org/abs/2407.08497.
In DF-QuAD, for any argument $\alpha\in\mathcal{A}$, $\sigma^{DF}(\alpha)$ is defined as follows:
Aggregation function:
$E_{\alpha}=\prod_{\left\{\beta\in\mathcal{A}\mid(\beta,\alpha)\in\mathcal{R^{-}}\right\}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!(1-\sigma^{DF}(\beta))-\prod_{\left\{\beta\in\mathcal{A}\mid(\beta,\alpha)\in\mathcal{R^{+}}\right\}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!(1-\sigma^{DF}(\beta)).$
Influence function:
$\sigma^{DF}(\alpha)=\begin{cases}\tau(\alpha)-\tau(\alpha)\cdot\left|E_{\alpha}\right|&if\ E_{\alpha}\leq 0;\\ \tau(\alpha)+(1-\tau(\alpha))\cdot E_{\alpha}&if\ E_{\alpha}>0.\\ \end{cases}$
An example of applying DF-QuAD is as follows.
Example 1.
Considering the QBAF in Figure 1, where $\tau(\alpha)=0.5$, $\tau(\beta)=0.3$, $\tau(\gamma)=0.6$, $\tau(\rho)=0.7$, and $\tau(\zeta)=0.4$. According to the aggregation and influence function of the DF-QuAD semantics [^3] We follow the convention that the product of an empty set is $1$ (not 00) because $1$ is the neutral element with respect to multiplication., we have $E_{\alpha}=0.24-0.91=-0.67$, $E_{\beta}=-0.7$, $E_{\gamma}=0.4$, $E_{\rho}=0$, $E_{\zeta}=0$, $\sigma^{DF}(\alpha)=\tau(\alpha)-\tau(\alpha)\cdot\left|E_{\alpha}\right|=0.165$, $\sigma^{DF}(\beta)=\tau(\beta)-\tau(\beta)\cdot\left|E_{\beta}\right|=0.09$, $\sigma^{DF}(\gamma)=\tau(\gamma)+(1-\tau(\gamma))\cdot E_{\gamma}=0.76$, $\sigma^{DF}(\rho)=\tau(\rho)=0.7$, $\sigma^{DF}(\zeta)=\tau(\zeta)=0.4$.
@@ 5.1 Algorithm Design Properties
We can restrict the update direction of arguments based on their polarity assuming that $\sigma$ respects directionality and monotonicity. Directionality (?) states that the strength of an argument depends only on its incoming edges.
@@ 5.2 Algorithms
Algorithm 1 computes the polarity from $\alpha$ to $\beta$ ($\alpha\neq\beta$) in three steps. Firstly, we compute all the non-cyclic paths from $\alpha$ to $\beta$ with Depth-First Search (DFS) and store them in a set of paths called $P_{\alpha\mapsto\beta}$. If there is no path, then the polarity is neutral. Secondly, all the nodes in all the paths are checked for cycles with the function find_cycles  [^5] See https://github.com/XiangYin2021/CE-QArg for details. A quicker implementation could involve applying (?), which is able to find all the elementary cycles of a directed graph in time bounded by O((n+e)(c+1)), where n, e, and c are the number of nodes, edges, and elementary cycles (nodes occur once except for the starting node) in the graph, respectively., which outputs a set of elementary cycles. If any of the node is part of a cycle, then we check whether each cycle contains an odd number of attacks. If this is the case, then the polarity from $\alpha$ to $\beta$ is unknown because the cycle will contain both odd and even numbers of attacks by going through the cycle an odd or even number of times. For instance, suppose a QBAF consists of two arguments $\alpha$ and $\beta$, where $\alpha$ attacks $\beta$, and $\beta$ supports $\alpha$. In this case, path $\langle(\alpha,\beta)\rangle$ is negative while $\langle(\alpha,\beta),(\beta,\alpha),(\alpha,\beta)\rangle$ is positive, thus $\alpha$ is unknown to $\beta$. Finally, we check the number of attacks on every path: if all the paths contain an odd (even) number of attacks, $\alpha$ is negative (positive) to $\beta$, and it is unknown otherwise.
@@
Algorithm 3 (which we call CE-QArg for Counterfactual Explanations for Quantitative bipolar Argumentation frameworks) is an iterative updating algorithm for identifying valid and cost-effective $\delta$-approximate counterfactuals. CE-QArg essentially involves determining the direction and magnitude of arguments. For brevity, we assume the current strength of the topic argument is less than the desired one. Firstly, we compute the polarity and priority for each argument with the function func_polarity and func_priority from Algorithm 1 and 2, respectively. Secondly, the key part of this algorithm is to determine the updating direction in every updating iteration. For this, we need an update list to record the updating direction in each iteration. For positive, negative and neutral arguments (lines 5-10), we only need to identify the updating direction once since they remain invariant, whereas, for unknown arguments (lines 13-20), we need to compute their updating direction in every iteration by difference quotient using the function func_dquo, which can be intuitively implemented by Definition 16 (so we omit the details). Once the updating direction of every argument is determined in an iteration, we update the base scores of all arguments all in one go by a small step multiplied by their priority (line 22) and make sure they are within the bounds. We assume the step is small enough that the interactions among arguments can be neglected. We repeat this procedure iteratively until the current strength reaches the desired strength, after which we return a possible counterfactual (lines 11-24).
@@ 7 Evaluations
We report on experiments with the QE semantics [^6] We give results with DF-QuAD and REB in https://arxiv.org/abs/2407.08497.. We set the updating step $\varepsilon$ to 0.01 in each iteration. Besides, we used both $L_{1}$ and $L_{2}$-norm distance as the metric for cost.
@@ 7.1 Experiment 1: Effectiveness
Table 1: Ablation studies for polarity and priority on acyclic/cyclic QBAFs: Comparison of average validity, $L_{1}$, $L_{2}$-norm distance, and runtime over 100 random generated acyclic full binary tree-like QBAFs with a depth of 7 and cyclic QBAFs with 100 arguments and 100 relations. [^8] We show the effectiveness on larger QBAFs in https://arxiv.org/abs/2407.08497.(The best results are shown in bold.)
@@
We first discuss the results of acyclic QBAFs (left side of slash in Table 8). All methods had the best validity of $1.00$. The use of priority is expected to shorten the $L_{1}$ and $L_{2}$-norm distance by updating the arguments close to the topic argument with a larger step. The results of BL and BL+pri showed that applying priority resulted in a 75.5% and 70.3% decrease in $L_{1}$ and $L_{2}$-norm distance, respectively, which also reduced the runtime by 23.7% in the meanwhile. These findings are in line with our hypothesis in that priority enables reaching a desired strength more cost-effectively, making the distance shorter and thus requiring fewer iterations. Utilizing polarity is expected to lower the runtime by computing the polarity only once for neutral, positive, and negative arguments in a QBAF. The results of BL and BL+pol showed that applying polarity significantly decreased the runtime by 98.7% as expected because the BL computes the difference quotient in every iteration for every argument, which causes the runtime wastage. For our CE-QArg, we see that the runtime is better than that of any of the previous three algorithms. Compared to BL+pri, although the $L_{1}$ and $L_{2}$-norm distance is slightly longer in our CE-QArg, the runtime is substantially decreased.
@@ 7.3 Experiment 3: Robustness
In addition, inspired by (??), we propose a robustness against noisy execution metric, which requires that the generated counterfactual could still lead to similar output (final strength for the topic argument) even if the counterfactual is perturbed. Still taking the loan application as an example, robustness against noisy execution ensures that when a rejected applicant is very close to the provided counterfactual, then the output of this counterfactual should also be close to the desired final strength. We evaluate robustness against noisy execution by the absolute difference between the strength of a topic argument obtained with two similar counterfactuals.
@@
Figure 7: Robustness evaluation of perturbing base score functions (left) and perturbing generated counterfactuals (right) over 100 random generated QBAFs with increasing perturbation $e$ from $10^{-8}$ to $10^{-1}$. Acyclic QBAFs are binary trees with a depth of 7 while cyclic QBAFs contain 100 arguments and relations. [^10] We show the robustness on larger QBAFs in https://arxiv.org/abs/2407.08497.
@@ 8 Conclusions
Acknowledgments
This research was partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101020934, ADIX) and by J.P. Morgan and by the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme. Any views or opinions expressed herein are solely those of the authors.
@@ Appendix D Additional Results for Section 7
Table 2: Ablation studies for polarity and priority on acyclic/cyclic QBAFs under the REB gradual semantics: Comparison of average validity, $L_{1}$, $L_{2}$-norm distance, and runtime over 100 random generated acyclic full binary tree-like QBAFs with a depth of 7 and cyclic QBAFs with 100 arguments and 100 relations. The best results are shown in bold.
Validity $L_{1}$ $L_{2}$ Runtime (s)
BL 0.99/0.77 13.89/18.35 1.45/1.93 2.34/77.59
BL+pri 1.00/0.93 3.02/0.44 0.39/0.27 0.94/28.77
BL+pol 1.00/1.00 22.32/1.50 2.08/0.48 0.02 /1.61
Ours 1.00/1.00 3.46/0.47 0.40/0.27 0.01/0.69
Table 3: Ablation studies for polarity and priority on acyclic QBAFs under the DF-QuAD gradual semantics: Comparison of average validity, $L_{1}$, $L_{2}$-norm distance, and runtime over 100 random generated acyclic full binary tree-like QBAFs with a depth of 7. The best results are shown in bold.
Validity $L_{1}$ $L_{2}$ Runtime (s)
BL 1.00 15.86 1.66 1.70
BL+pri 1.00 6.89 0.85 2.67
BL+pol 1.00 25.81 2.43 0.02
Ours 1.00 9.47 1.02 0.02
Figure 8: Scalability evaluation for CE-QArg on acyclic (left) and cyclic (right) QBAFs under REB gradual semantics: comparison of average runtime over 100 randomly generated acyclic and cyclic QBAFs.
Figure 9: Scalability evaluation for CE-QArg on acyclic QBAFs under DF-QuAD gradual semantics: comparison of average runtime over 100 randomly generated acyclic QBAFs.
Figure 10: Robustness evaluation of perturbing base score functions (left) and perturbing generated counterfactuals (right) over 100 random generated QBAFs with increasing perturbation $e$ from $10^{-8}$ to $10^{-1}$ under REB gradual semantics. Acyclic QBAFs are binary trees with a depth of 7 while cyclic QBAFs contain 100 arguments and relations.
Figure 11: Robustness evaluation of perturbing base score functions (left) and perturbing generated counterfactuals (right) over 100 random generated QBAFs with increasing perturbation $e$ from $10^{-8}$ to $10^{-1}$ under DF-QuAD gradual semantics. Acyclic QBAFs are binary trees with a depth of 7.